A 50 g slug slides down a 2.0 m long drain pipe. If the slug's speed at the bottom of the pipe is only 3.0 m/s, how much energy was dissipated as heat caused by friction?


1
Expert's answer
2021-01-13T11:35:55-0500
"Q=mgh-0.5mv^2\\Q=(0.05)((9.8)(2)-0.5(3)^2)=0.755\ J"